OpenStudy (anonymous):

How many pounds of nuts worth $1.05 a pound must be mixed with nuts worth $0.85 a pound to get a mixture of 200 pounds of nuts worth $0.90 a pound?

OpenStudy (pfenn1):

If x=number of pounds of nut worth $1.05/pound and y=number of pounds of nuts worth $0.85/pound, we know two things. x+y=200 1.05x + 0.85y =200 (0.9) solve equation 1 for x=200-y, substitute into the 2nd equation and you have your answer.